### v1

@@ 3.5 Optimization
4 The ArTVideo Benchmark
4.1 Data Source and Annotation
Data Annotation. As shown in Fig. 4, we provide high-quality word-level annotations for both straight and curved text in two different annotated styles. Straight text is labeled with quadrilaterals, while for curved text, we follow the CTW1500 [21] and adopt a polygon with 14 points to annotate the text contour. To record the details of each text instance, we annotate the following attributes: 1) ID refers to the unique identifier assigned to each text instance across consecutive frames; 2) Text Category is classified into two types: ‘alphanumeric’, which includes Latin text, and ‘other’, which encompasses both blurry and non-Latin text; 3) Transcription means the text string for each text region; 4) Bbox and Poly represent the bounding boxes of text instances, in horizontal and polygonal formats, respectively; 5) Segmentation denotes the mask annotation for each text region; 6) Box Type indicates the type of box annotation for each text instance, which is either ‘quadrilateral’ (for straight text) or ‘polygon’ (for curved text).
@@ 4.1 Data Source and Annotation
4.2 Statistical Analysis
Dataset Characteristics. The details of ArTVideo, along with a comparison to other existing video text datasets, are presented in Tab. 1. The table clearly highlights three key characteristics of ArTVideo: 1) It contains over 34% curved text, whereas other datasets primarily consist of straight text; 2) The annotation boxes are in polygonal form, enhancing annotation precision; 3) ArTVideo includes additional mask annotations, allowing models trained on this dataset to produce more precise segmentation results for text regions. These distinctive characteristics make ArTVideo a valuable resource for advancing video text spotting research, particularly in scenarios involving curved text and the need for precise segmentation results.
@@ 4.2 Statistical Analysis
Word Cloud. We present a word cloud visualization of the textual content in Fig. 5 (c). The most frequently occurring word in this dataset is ‘THE’ followed by ‘OF’, ‘ON’, ‘HOT’, ‘HER’, ‘GRACE’, and other words with lengths of no more than 10 characters. These observations are consistent with the statistical results of the text length distribution shown in panel (b).
4.3 Supported Tasks and Metrics
@@
Following ICDAR15-video [14], BOVText [33] and DSText [34], the tracking and spotting performance of the models are assessed with three metrics: MOTA (Multiple Object Tracking Accuracy) [3], MOTP (Multiple Object Tracking Precision), and IDF1 [26]. MOTA indicates the overall performance of detection and tracking but prefers to reflect detection performance. MOTP measures the average spatial distance between the predicted locations of tracked objects and the ground truth positions. IDF1 provides a measurement of association accuracy.
5 Experiments on Open-source Datasets
5.1 Datasets and Evaluation Metrics
ICDAR13-video [16] contains 13 videos for training and 15 videos for testing from indoors and outdoors scenarios with quadrilateral annotations at word-level.
ICDAR15-video [14] is an extended version of ICDAR2013-video, serving as a word-level video text benchmark with quadrilateral annotations, comprising 25 training and 24 test videos, focusing on wild scenarios.
BOVText [33] is a bilingual video text spotting benchmark (English and Chinese) with textline-level quadrilateral annotations, covering different scenarios in 2,021 videos, such as ‘Cartoon’, ‘Vlog’, ‘Movie’, and more.
DSText [34] provides 50 training videos and 50 test videos. Compared with the previous datasets, DSText mainly includes the following three new challenges: dense video texts, high-proportioned small texts, and various new scenarios, e.g., ‘Game’, ‘Sports’, etc. Similar to ICDAR15-video, DSText adopts word-level annotations, which are labeled with quadrilaterals.
@@ 5.1 Datasets and Evaluation Metrics
5.2 Implementation Details
In all experiments, we only use a single NVIDIA GeForce RTX 3090 (24G) GPU to train and evaluate GoMatching and GoMatching++. As for the image text spotter, we apply the officially released DeepSolo [37]. To adapt GoMatching and GoMatching++ for downstream video datasets, we only update the rescoring head and LST-Matcher, while keeping DeepSolo frozen.
Table 2: Impact of difference components in the proposed GoMatching. ‘Query’ indicates that LST-Matcher employs the queries of high-score text instances for association, otherwise RoI features. ‘Scoring’ indicates the employed scoring mechanism, in which ‘O’ means using the original scores from DeepSolo, ‘R’ means using the scores recomputed by the rescoring head, and ‘F’ means using the fusion scores obtained from the rescoring mechanism.
Index Query Scoring LT-Matcher ST-Matcher MOTA ($\uparrow$) MOTP ($\uparrow$) IDF1 ($\uparrow$)
1 O $\checkmark$ 66.20 78.52 75.07
2 $\checkmark$ O $\checkmark$ 67.22 78.54 76.12
3 $\checkmark$ R $\checkmark$ 68.47 78.29 77.09
4 $\checkmark$ F $\checkmark$ 68.80 78.24 77.41
5 $\checkmark$ F $\checkmark$ 69.40 78.34 73.60
6 $\checkmark$ F $\checkmark$ $\checkmark$ 70.52 78.25 78.70
@@ 5.2 Implementation Details
Table 3: Video text detection performance on ICDAR2013-video. The best results are highlighted in bold.
Method Prec. ($\uparrow$) Rec. ($\uparrow$) F-measure ($\uparrow$)
Free [5] 79.7 68.4 73.6
TransDETR [29] 80.6 70.2 75.0
GoMatching w/o rescoring 92.4 65.7 76.8
GoMatching 89.5 74.8 81.5
Training Setting. The image text spotting part is initialized with off-the-shelf DeepSolo weights and remains frozen in all experiments. We optimize other modules on video datasets. We follow EfficientDet [28] to adopt the scale-and-crop augmentation strategy with a resolution of 1,280. The batch size $B$ is 6. All frames in a batch are from the same video. AdamW [19] is used as the optimizer. We adopt the warmup cosine annealing learning rate strategy with the initial learning rate being set to $5\times 10^{-5}$. The loss weights $\lambda_{res}$ and $\lambda_{asso}$ are set to 1.0 and 0.5, respectively. For focal loss, $\alpha$ is 0.25 and $\gamma$ is 2.0 as in [6, 37]. The model is trained for 30k iterations on all downstream video datasets.
5.3 Ablation Studies
Effectiveness of Utilizing Queries. A comparison of the first two rows in Tab. 2 reveals that using queries from the decoder of image text spotter is more beneficial for tracking than RoI features [29]. Leveraging queries from frozen DeepSolo achieves 0.98% and 1.05% improvements on MOTA and IDF1. This improvement can be attributed to the fact that queries integrate more text instance information, i.e., unifying multi-scale features, text semantics, and position information, which has been proven effective in DeepSolo. Although position information is essential for tracking, it is ignored in RoI features.
Effectiveness of Rescoring Mechanism. To verify the effectiveness of the rescoring mechanism, we test three different scoring mechanisms: the original scores from DeepSolo, the scores recomputed by the rescoring head, and the fusion scores from the rescoring mechanism. As presented in row 2 and row 3 of Tab. 2, the rescoring head effectively alleviates the performance degradation caused by the domain gap between ICDAR15-image and ICDAR15-video and achieves gains of 1.25% and 0.97% on MOTA and IDF1, respectively. Moreover, as shown in row 4, combining the knowledge of rescoring head learned from the video dataset with the prior knowledge of DeepSolo can further improve MOTA and IDF1 by 0.33% and 0.32%, respectively.
Following previous works [5, 29], we evaluate the performance of GoMatching for video text detection on ICDAR13-video. As shown in Tab. 3, without the rescoring mechanism, GoMatching depends on the original outputs from DeepSolo, resulting in a 4.5% decrease in Recall and only a 1.8% improvement in F-measure compared to TransDETR. This performance drop is due to the domain gap between image and video data, as directly applying an image text spotter yields low-confidence results and poor Recall on video data. However, with the rescoring mechanism, GoMatching achieves a 9.1% improvement in Recall over DeepSolo and a 6.5% boost in F-measure over TransDETR. These gains highlight the effectiveness of the rescoring mechanism in mitigating the domain gap and improving the tracking candidate pool, demonstrating that GoMatching’s strong performance is not solely due to the use of a robust image text spotter.
Table 4: Comparison results of detection AP on the ICDAR13-video with and without the rescoring mechanism.
Method AP $\text{AP}_{\text{S}}$ $\text{AP}_{\text{M}}$ $\text{AP}_{\text{L}}$
w/o rescoring 26.2 11.6 40.1 49.8
w/ rescoring 29.3 (+3.1) 15.5 (+3.9) 42.7 (+2.6) 51.9 (+2.1)
To further explore how the rescoring mechanism eases the domain gap, we calculate the AP of detection results on ICDAR13-video. As shown in Tab. 4, incorporating the rescoring mechanism effectively improves the detection performance of DeepSolo on video datasets, particularly for small text, resulting in a 3.9% improvement on $\text{AP}_{\text{S}}$.
Table 5: Results of different score fusion strategies on ICDAR5-video. ‘Mean’, ‘Geo-mean’, and ‘Maximum’ denote the arithmetic mean, geometric mean, and the maximum score fusion strategies, respectively. The best results are highlighted in bold.
Strategy MOTA ($\uparrow$) MOTP ($\uparrow$) IDF1 ($\uparrow$)
Mean 70.46 78.38 78.29
Geo-mean 70.29 78.39 78.26
Maximum 70.52 78.25 78.70
Figure 6: Visualization of ST-Matcher and LST-Matcher. The first row illustrates a failure case with ID switches due to missed detections and incorrect matching using only ST-Matcher. The second row demonstrates how LST-Matcher effectively addresses this issue by combining both long- and short-term matching. Text instances in the same color correspond to the same IDs.
Different Score Fusion Strategies. To investigate the impact of the score fusion strategy in the rescoring mechanism, we further evaluate two alternatives: (1) the arithmetic mean strategy (mean), and (2) the geometric mean strategy (geo-mean). The arithmetic mean strategy takes the average of the scores from the image text spotter and the rescoring head as the final score for each query, while the geo-mean strategy uses the geometric mean. These two strategies can be formulated as:
$c_{mean}=(c_{o}+c_{r})/2,$ (10)
$c_{geo-mean}=\sqrt{c_{o}*c_{r}},$ (11)
where $c_{mean}$ and $c_{geo-mean}$ denote the final score of the two strategies, respectively. $c_{o}$ and $c_{r}$ are the scores from the image text spotter and the rescoring head, respectively.
@@ 5.3 Ablation Studies
Table 6: Video text spotting results for different LST-Matcher architectural designs. ‘T-Para’ denotes the number of trainable parameters.
@@
Table 7: Comparison with SOTA methods on three open-source datasets. ‘†’ indicates results from the official competition website, and ‘*’ denotes results from the officially released model. ‘M-ME’ shows whether multi-model ensembling was used (‘Y’ for Yes, ‘N’ for No). The best and second-best results are marked in bold and underlined, respectively.
@@
ICDAR15-video HIK_OCR [5] 43.16 76.78 57.92 52.98 74.88 61.85 N
CoText [31] 50.3 73.6 67.4 58.94 74.53 71.66 N
TransDETR [29] 48.86 74.01 65.33 60.96 74.61 72.80 N
h&h_lab† 53.74 76.92 65.51 63.76 77.78 71.08 Y
GOCR Offline† 54.54 73.61 71.21 63.90 74.46 77.56 -
@@
BOVText PSENet (+ RARE) [31] 52.1 77.5 45.9 -16.2 80.3 31.7 Y
DB (+ RARE) [31] 53.2 78.3 48.3 -12.6 81.1 34.2 Y
TransVTSpotter [31] 68.2 82.1 64.7 -1.4 82.0 40.8 N
CoText [31] 69.7 80.3 76.5 10.3 80.4 47.1 N
@@
As shown in Tab. 5, the maximum score fusion strategy outperforms all other strategies on MOTA and IDF1. In contrast, the other strategies, which rely on low-confidence scores from the image text spotter, tend to produce lower final scores and may lead to missed detections. Therefore, we use the maximum score fusion strategy by default in GoMatching’s rescoring mechanism.
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 7: Visual results of video text spotting. From top to bottom: results on ICDAR15-video, BOVText, and DSText. Text instances associated with the same trajectories are assigned the same colors.
Effectiveness of LST-Matcher. In this part, we conduct three experiments to evaluate the effectiveness of the LST-Matcher. As shown in row 4 of Tab. 2, LT-Matcher is used to associate text instances in the current frame with trajectories stored in the tracking memory bank. Row 5 shows the use of ST-Matcher, which associates text instances in the current frame with trajectories from the previous frame. Finally, row 6 presents the results of combining both LT-Matcher and ST-Matcher in the LST-Matcher. We observe that LST-Matcher outperforms LT-Matcher by 1.72% in MOTA and 1.29% in IDF1, and surpasses ST-Matcher by 1.12% in MOTA and 5.1% in IDF1. Additionally, as shown in Fig. 6, LST-Matcher effectively reduces ID switches caused by missed detections or significant appearance changes due to motion blur. These results demonstrate that combining short-term and long-term information improves tracking robustness, thereby enhancing video text spotting performance.
Different Architectures of LST-Matcher. We investigate optimal architectural designs for LST-Matcher, with results summarized in Tab. 6. For fairly performance comparison across variants, we maintain consistent structure and parameters for all other components. Tab. 6 shows that removing either FFN, which maps queries to a more suitable feature space, or the attention mechanism, which focuses on key instance features, leads to a significant drop in performance. Notably, a simplified design, which reducing the attention mechanism to cross-attention only, achieves a two-thirds reduction in parameters while maintaining or even surpassing the performance of transformer-based architectures across all datasets. This simple yet effective design highlights the strong potential of GoMatching++ as a parameter-efficient baseline for video text spotting.
Table 8: Results of using different image sizes on ICDAR15-video. The values in parentheses mean the size of the shorter side of the input image during inference. The best results are highlighted in bold.
@@
Table 9: Comparison between TransDETR and GoMatching. ‘T-Para.’, ‘Mem.’, and ‘GPU-hours’ refer to the number of trainable parameters, memory usage, and GPU hours consumed during training, respectively.
@@
5.4 Comparison with State-of-the-art Methods
Results on ICDAR15-video. To evaluate the effectiveness of our method on oriented video text, we compare it with SOTA approaches on ICDAR15-video, as shown in Tab. 7. GoMatching and GoMatching++ outperform all other methods across all evaluation metrics, with GoMatching++ achieving the best performance. By integrating a robust image text spotter with a powerful tracker, GoMatching++ improves performance by 7.87% MOTA, 1.02% MOTP, and 1.66% IDF1 for tracking, and by 5.24% MOTA, 0.74% MOTP, and 2.55% IDF1 for spotting. Additionally, GoMatching++ surpasses the current SOTA single-model method, TransDETR, with improvements of 13.55% MOTA, 3.93% MOTP, and 7.54% IDF1 for tracking, and 11.08% MOTA, 3.92% MOTP, and 7.31% IDF1 for spotting.
Results on BOVText. Our methods, except for English text, are easily adaptable to other video text scenarios, such as bilingual textline spotting. For BOVText, which focuses on English and Chinese textline spotting, we use DeepSolo trained on bilingual datasets and fine-tune GoMatching and GoMatching++ on BOVText. The results, shown in Tab. 7, demonstrate that GoMatching++ sets a new record, outperforming previous methods significantly, with the exception of the IDF1 metric in video tracking. Notably, GoMatching++ excels in video text spotting, surpassing the previous SOTA method, CoText [31], with improvements of 42.6% in MOTA, 6.8% in MOTP, and 15.7% in IDF1. This outstanding performance on BOVText highlights GoMatching++’s ability to effectively spot both Chinese and English text in videos, and its flexibility for other languages through customizable image text spotters. The gap in tracking and spotting results of prior methods further underscores that recognition ability remains the key limitation.

### v2

@@ IV-A Data Source and Annotation
Data Annotation. As shown in Fig. 4, we provide high-quality word-level annotations for both straight and curved text in two different annotated styles. Straight text is labeled with quadrilaterals, while for curved text, we follow the CTW1500 [liu2019curved] and adopt a polygon with 14 points to annotate the text contour. To record the details of each text instance, we annotate the following attributes: 1) ID refers to the unique identifier assigned to each text instance across consecutive frames; 2) Text Category is classified into two types: ‘alphanumeric’, which includes Latin text, and ‘other’, which encompasses both blurry and non-Latin text; 3) Transcription means the text string for each text region; 4) Bbox and Poly represent the bounding boxes of text instances, in horizontal and polygonal formats, respectively; 5) Segmentation denotes the mask annotation for each text region; 6) Box Type indicates the type of box annotation for each text instance, which is either ‘quadrilateral’ (for straight text) or ‘polygon’ (for curved text).
@@
IV-B Statistical Analysis
Dataset Characteristics. The details of ArTVideo, along with a comparison to other existing video text datasets, are presented in Tab. I. The table clearly highlights three key characteristics of ArTVideo: 1) It contains over 34% curved text, whereas other datasets primarily consist of straight text; 2) The annotation boxes are in polygonal form, enhancing annotation precision; 3) ArTVideo includes additional mask annotations, allowing models trained on this dataset to produce more precise segmentation results for text regions. These distinctive characteristics make ArTVideo a valuable resource for advancing video text spotting research, particularly in scenarios involving curved text and the need for precise segmentation results.
@@ IV-B Statistical Analysis
Word Cloud. We present a word cloud visualization of the textual content in Fig. 5 (c). The most frequently occurring word in this dataset is ‘THE’ followed by ‘OF’, ‘ON’, ‘HOT’, ‘HER’, and other words with lengths of no more than 10 characters. These observations are consistent with the statistical results of the text length distribution shown in panel (b).
IV-C Supported Tasks and Metrics
@@
Following ICDAR15-video [Karatzas2015ICDAR15], BOVText [wu2021bilingual] and DSText [Wu2023DSText], the tracking and spotting performance of the models are assessed with three metrics: MOTA (Multiple Object Tracking Accuracy) [bernardin2008evaluating], MOTP (Multiple Object Tracking Precision), and IDF1 [ristani2016performance]. MOTA indicates the overall performance of detection and tracking but prefers to reflect detection performance. MOTP measures the average spatial distance between the predicted locations of tracked objects and the ground truth positions. IDF1 provides a measurement of association accuracy.
V Experiments on Open-source Datasets
V-A Datasets and Evaluation Metrics
ICDAR13-video [karatzas2013icdar] contains 13 videos for training and 15 videos for testing from indoors and outdoors scenarios with quadrilateral annotations at word-level.
ICDAR15-video [Karatzas2015ICDAR15] is an extended version of ICDAR2013-video, serving as a word-level video text benchmark with quadrilateral annotations, comprising 25 training and 24 test videos, focusing on wild scenarios.
BOVText [wu2021bilingual] is a bilingual video text spotting benchmark (English and Chinese) with textline-level quadrilateral annotations, covering different scenarios in 2,021 videos, such as ‘Cartoon’, ‘Vlog’, ‘Movie’, and more.
DSText [Wu2023DSText] provides 50 training videos and 50 test videos. Compared with the previous datasets, DSText mainly includes the following three new challenges: dense video texts, high-proportioned small texts, and various new scenarios. Similar to ICDAR15-video, DSText adopts word-level annotations, which are labeled with quadrilaterals.
@@ V-A Datasets and Evaluation Metrics
TABLE II: Impact of difference components in the proposed GoMatching++.
Index ReScoring Matcher MOTA ($\uparrow$) MOTP ($\uparrow$) IDF1 ($\uparrow$)
1 LST-Matcher 70.68 78.73 78.42
2 $\checkmark$ ST-Matcher 71.17 78.41 75.53
3 $\checkmark$ LT-Matcher 70.07 78.52 78.48
4 $\checkmark$ LST-Matcher 72.20 78.52 80.11
@@
TABLE III: Video text detection performance on ICDAR2013-video. The best results are highlighted in bold.
Method Prec. ($\uparrow$) Rec. ($\uparrow$) F-Score ($\uparrow$)
Free [cheng2020free] 79.7 68.4 73.6
TransDETR [wu2024end] 80.6 70.2 75.0
GoMatching++ w/o rescoring 92.4 65.7 76.8
GoMatching++ 89.5 74.8 81.5
@@
TABLE IV: Video text spotting results for different LST-Matcher architectural designs. ‘T-Para’ denotes the number of trainable parameters.
@@
V-B Implementation Details
In all experiments, we only use a single NVIDIA GeForce RTX 3090 (24G) GPU to train and evaluate GoMatching and GoMatching++. As for the image text spotter, we apply the officially released DeepSolo [ye2023deepsolo]. To adapt GoMatching and GoMatching++ for downstream video datasets, we only update the rescoring head and LST-Matcher, while keeping DeepSolo frozen.
Training Setting. The image text spotting part is initialized with off-the-shelf DeepSolo weights and remains frozen in all experiments. We optimize other modules on video datasets. We follow EfficientDet [tan2020efficientdet] to adopt the scale-and-crop augmentation strategy with a resolution of 1,280. The batch size $B$ is 6. All frames in a batch are from the same video. AdamW [loshchilov2018decoupled] is used as the optimizer. We adopt the warmup cosine annealing learning rate strategy with the initial learning rate being set to $5\times 10^{-5}$. The loss weights $\lambda_{res}$ and $\lambda_{asso}$ are set to 1.0 and 0.5, respectively. For focal loss, $\alpha$ is 0.25 and $\gamma$ is 2.0 as in [carion2020end, ye2023deepsolo]. The model is trained for 30k iterations on all downstream video datasets.
V-C Ablation Studies
Effectiveness of Rescoring Mechanism. As presented in row 1 and row 4 of Tab. III, the rescoring mechanism can effectively combining the knowledge of rescoring head learned from the video dataset with the prior knowledge of DeepSolo, resulting performance gains of 1.52% and 1.69% on MOTA and IDF1, respectively. Moreover, following previous works [cheng2020free, wu2024end], we evaluate the performance of GoMatching++ for video text detection on ICDAR13-video. Since GoMatching and GoMatching++ employ the same text detector, we only report the video text detection performance of GoMatching++. As shown in Tab. III, without the rescoring mechanism, GoMatching++ depends on the original outputs from DeepSolo, resulting in a 4.5% decrease in recall and only a 1.8% improvement in F-Score compared to TransDETR. This performance drop is due to the domain gap between image and video data, as directly applying an image text spotter yields low-confidence results and poor recall on video data. However, with the rescoring mechanism, GoMatching++ achieves a 9.1% improvement in recall over DeepSolo and a 6.5% boost in F-Score over TransDETR. These gains highlight the effectiveness of the rescoring mechanism in mitigating the domain gap and improving the tracking candidate pool, demonstrating that GoMatching++’s strong performance is not solely due to the use of a robust image text spotter.
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 6: Visual results of video text spotting achieved by GoMatching++. From top to bottom: results on ICDAR15-video, BOVText, and DSText. Text instances associated with the same trajectories are assigned the same colors.
TABLE V: Comparison with SOTA methods on three open-source datasets. ‘†’ indicates results from the official competition website, and ‘*’ denotes results from the officially released model. ‘M-ME’ shows whether multi-model ensembling was used (‘Y’ for Yes, ‘N’ for No). The best and second-best results are marked in bold and underlined, respectively.
@@ V-C Ablation Studies
ICDAR15-video HIK_OCR [cheng2020free] 43.16 76.78 57.92 52.98 74.88 61.85 N
CoText [wu2024bilingual] 50.3 73.6 67.4 58.94 74.53 71.66 N
TransDETR [wu2024end] 48.86 74.01 65.33 60.96 74.61 72.80 N
@@
BOVText TransVTSpotter [wu2021bilingual] 68.2 82.1 64.7 -1.4 82.0 40.8 N
CoText [wu2024bilingual] 69.7 80.3 76.5 10.3 80.4 47.1 N
@@
Effectiveness of LST-Matcher. In this part, we conduct three experimental settings to evaluate the effectiveness of the LST-Matcher. ST-Matcher associates text instances in the current frame with trajectories from the previous frame, while LT-Matcher associates text instances in the current frame with trajectories stored in the tracking memory bank. The LST-Matcher combines both ST-Matcher and LT-Matcher. As shown in rows 2–4 of Tab. III, LST-Matcher outperforms ST-Matcher by 1.03% in MOTA and 4.58% in IDF1, and surpasses LT-Matcher by 2.13% in MOTA and 1.63% in IDF1. The results demonstrate that integrating short-term and long-term information can improves tracking robustness, thereby enhancing video text spotting performance.
Different Architectures of LST-Matcher. We investigate optimal architectural designs for LST-Matcher, with results summarized in Tab. IV. For fairly performance comparison across variants, we maintain consistent structure and parameters for all other components. Tab. IV shows that removing either FFN, which maps queries to a more suitable feature space, or the attention mechanism, which focuses on key instance features, leads to a significant drop in performance. Notably, a simplified design, which reducing the attention mechanism to cross-attention only, achieves a two-thirds reduction in parameters while maintaining or even surpassing the performance of transformer-based architectures across all datasets. This simple yet effective design highlights the strong potential of GoMatching++ as a parameter-efficient baseline for video text spotting.
TABLE VI: Results of using different image sizes on ICDAR15-video. The values in parentheses mean the size of the shorter side of the input image during inference. The best results are highlighted in bold.
@@
TABLE VII: Comparison between TransDETR and GoMatching. ‘T-Para.’, ‘Mem.’, and ‘GPU-hours’ refer to the number of trainable parameters, memory usage, and GPU hours consumed during training, respectively.
@@
V-D Comparison with State-of-the-art Methods
Results on ICDAR15-video. To evaluate the effectiveness of our method on oriented video text, we compare it with SOTA approaches on ICDAR15-video, as shown in Tab. V. GoMatching and GoMatching++ outperform all other methods across all evaluation metrics, with GoMatching++ achieving the best performance. By integrating a robust image text spotter with a powerful tracker, GoMatching++ improves performance by 7.87% MOTA, 1.02% MOTP, and 1.66% IDF1 for tracking, and by 5.24% MOTA, 0.74% MOTP, and 2.55% IDF1 for spotting. Additionally, GoMatching++ surpasses the current SOTA single-model method, TransDETR, with improvements of 13.55% MOTA, 3.93% MOTP, and 7.54% IDF1 for tracking, and 11.08% MOTA, 3.92% MOTP, and 7.31% IDF1 for spotting.
Results on BOVText. Our methods, except for English text, are easily adaptable to other video text scenarios, such as bilingual textline spotting. For BOVText, which focuses on English and Chinese textline spotting, we use DeepSolo trained on bilingual datasets and fine-tune GoMatching and GoMatching++ on BOVText. The results, shown in Tab. V, demonstrate that GoMatching++ sets a new record, outperforming previous methods significantly, with the exception of the IDF1 metric in video tracking. Notably, GoMatching++ excels in video text spotting, surpassing the previous SOTA method, CoText [wu2024bilingual], with improvements of 42.6% in MOTA, 6.8% in MOTP, and 15.7% in IDF1. This outstanding performance on BOVText highlights GoMatching++’s ability to effectively spot both Chinese and English text in videos, and its flexibility for other languages through customizable image text spotters. The gap in tracking and spotting results of prior methods further underscores that recognition ability remains the key limitation.
Results on DSText. We further conducted experiments on DSText, focusing on dense and small video text scenarios, with results shown in Tab. V. Notably, many top methods on the DSText leaderboard rely on model ensembles and large public datasets to boost performance [Wu2023DSText]. For instance, TencentOCR combines detection outputs from DBNet [liao2020real] and Cascade MaskRCNN [cai2019cascade], integrates the Parseq [bautista2022scene] text recognizer, and improves end-to-end tracking with ByteTrack [zhang2022bytetrack]. Similarly, DA uses Mask R-CNN [he2017mask] and DBNet for text detection, BotSORT [aharon2022bot] to replace the tracker in VideoTextSCM [gao2021video], and Parseq for recognition. As a single model, GoMatching++ slightly trails behind ensemble-based methods in video text tracking. However, this performance gap can be reduced with stronger detectors. On video text spotting task, our method shows strong capabilities, ranking first in MOTA (23.23%), second in MOTP (80.43%), and third in IDF1 (46.24%). Compared to the best single-model approach, GoMatching++ achieves significant improvements: +9.04% in MOTA and +2.83% in IDF1 for video text tracking, and +45.86% in MOTA and +19.81% in IDF1 for video text spotting.